# PRIMER: Perception-Aware Robust Learning-based Multiagent Trajectory Planner

### v1

@@ Abstract
In decentralized multiagent trajectory planners, agents need to communicate and exchange their positions to generate collision-free trajectories. However, due to localization errors/uncertainties, trajectory deconfliction can fail even if trajectories are perfectly shared between agents. To address this issue, we first present PARM and PARM*, perception-aware, decentralized, asynchronous multiagent trajectory planners that enable a team of agents to navigate uncertain environments while deconflicting trajectories and avoiding obstacles using perception information. PARM* differs from PARM as it is less conservative, using more computation to find closer-to-optimal solutions. While these methods achieve state-of-the-art performance, they suffer from high computational costs as they need to solve large optimization problems onboard, making it difficult for agents to replan at high rates. To overcome this challenge, we present our second key contribution, PRIMER, a learning-based planner trained with imitation learning (IL) using PARM* as the expert demonstrator. PRIMER leverages the low computational requirements at deployment of neural networks and achieves a computation speed up to $5500$ times faster than optimization-based approaches.
@@ I Introduction
In recent years, multiagent UAV trajectory planning has been extensively studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In real-world deployments of multiagent trajectory planning methods, it is crucial to deal with challenges such as (1) detecting and avoiding collisions with unknown obstacles, (2) handling localization errors/uncertainties, (3) achieving scalability to a large number of agents, and (4) enabling fast and efficient computation for onboard replanning and quick adaptation to dynamic environments. However, finding effective solutions to these challenges remains an open question.
@@
Many optimization-based approaches [4, 5, 12, 6] have been proposed for multiagent trajectory generation. However, these approaches often require substantial computational resources, posing challenges for deployments in dynamic environments that demand fast on-the-fly replanning. To mitigate this issue, researchers have explored imitation learning (IL)-based approaches [16, 17, 18], which offer the advantage of faster replanning while still achieving close-to-optimal trajectory generation.
TABLE I: State-of-the-art UAV Trajectory Planners
Method Multiagent Perception-aware
EGO-Swarm [6]
DMPC [19]
MADER [4] Yes No
decMPC [20]
RMADER [5]
Raptor [21]
Time-opt [22]
PANTHER [23] No Yes
PA-RHP [24]
Deep-PANTHER [16]
PARM / PRIMER (proposed) Yes Yes
@@
    PARM/PARM* – the first decentralized, asynchronous, perception-aware, multiagent trajectory planner.
@@ II Perception-aware Multiagent Trajectory Generation
This section outlines our perception-aware multiagent planning approach, encompassing both optimization-based and IL-based methods. First, we introduce PARM and PARM*, which are optimization-based perception-aware multiagent trajectory planners. These planners utilize the Robust MADER trajectory deconfliction framework [5] to ensure safe and collision-free trajectories. By considering both position and yaw, PARM and PARM* are capable of tracking multiple obstacles while generating trajectories. Next, we present PRIMER, an IL-based planner that is trained using PARM* as an expert. PRIMER offers fast computation by leveraging the benefits of IL. By incorporating the expertise of PARM*, PRIMER is able to generate effective trajectories while significantly reducing computational requirements. As Table II shows, PARM, PARM*, and PRIMER are notable as the first perception-aware multiagent trajectory planners designed to track multiple obstacles while accounting for both position and yaw.
@@ II-A PARM / PARM* —Optimization-based Decentralized, Asynchronous Perception-Aware Multiagent Planning
MADER [4] proposed an optimization-check-recheck scheme for decentralized, asynchronous multiagent planning. In this approach, an agent optimizes its trajectory while using received trajectories as optimization constraints. Next, the agent checks its trajectory against trajectories received in the optimization step and rechecks if it received any trajectory in the check step. To enhance robustness against communication delays, we proposed Robust MADER [5], which replaces the recheck step with a delay-check step. These frameworks allow fully decentralized asynchronous multiagent trajectory generation under real-world uncertainties and delays.
PANTHER [23] proposed a perception-aware trajectory planner for a single agent in dynamic environments, generating trajectories to avoid obstacles while keeping them in the sensor FOV. In [16], PANTHER* improved the original PANTHER with less conservatism by including separating planes and trajectory total time as optimization variables, but both were limited to tracking and avoiding only one obstacle at a time. To overcome this limitation, we modify the optimization problem solved by PANTHER to enable tracking and avoidance of multiple obstacles, leading to PARM – a decentralized, asynchronous, perception-aware multi-agent trajectory planning system that incorporates this modified optimization approach into the RMADER deconfliction framework. To enable tracking multiple obstacles and agents, we modify the FOV term given in [23, Section IV] as the following.
@@
Additionally, following the method outlined in [16], we introduce PARM*, a method that finds closer-to-optimal trajectories, albeit at the cost of increased computation and slower re-planning rates.
@@
In addition, it is worth noting that coupled trajectory generation has the advantage of enabling the optimizer to simultaneously consider both position and yaw trajectories, taking into account the impact of each on the other. On the other hand, the decoupled approach involves optimizing either the position or yaw trajectory first, and then optimizing the other trajectory based on the pre-determined trajectory. While this approach may reduce computation and complexity, it could also result in sub-optimal position and yaw trajectories overall, as the two trajectories are not jointly optimized. Therefore we designed PARM/PARM* in a way that they optimize both position and yaw in a coupled manner (See Table II).
@@ III Simulation Results
Table III compares the average performance of PRIMER and PARM* in a simulation environment with a single dynamic obstacle that follows a trefoil trajectory while the agent flies diagonally to avoid obstacles. The comparison is based on the average cost of trajectories, computation time, obstacle avoidance failure rate, and dynamic constraint failure rate.
@@ III-B Multiagent and multi-obstacle benchmarking
Table IV compares the average performance of PARM, PARM*, and PRIMER on two different environments: (1) one agent with two obstacles, and (2) three agents with two obstacles. The metrics used to evaluate the performance are as follows:
@@
    Computation time: the time it takes to replan at each step.
@@
    Translational dynamic constraints violation rate: the violation rate of the maximum velocity, acceleration, and jerk.
@@
    Yaw dynamic constraints violation rate: the violation rate of the maximum yaw rate.
@@
Fig. 5 illustrates (a) Computation Time, (b) Travel Time, and (c-d) Trajectory Smoothness. While the 6-trajectory approach finds smoother trajectories, it requires much longer computation time, resulting in increased travel time. The trajectories generated by PRIMER exhibit smoothness comparable to those of PARM*, while PARM and PRIMER have significantly faster computation times compared to PARM*.
@@ IV Conclusions
To overcome this challenge, we presented PRIMER, a learning-based planner that was trained with imitation learning (IL) using PARM* as the expert demonstrator. PRIMER is a computationally-efficient deep neural network and achieved a computation speedup of up to $5500$ times faster than optimization-based approaches, while maintaining high performance. This speedup enables scalability to a large number of agents, making PRIMER a promising approach for large-scale swarm coordination.
Moving forward, our future work will focus on larger-scale simulations and hardware flight experiments to demonstrate the scalability and performance of PRIMER in complex environments with many agents and obstacles. Additionally, we will explore how to integrate PRIMER with other state-of-the-art perception systems, such as SLAM, to enable even more robust and accurate perception-aware multiagent trajectory planning. Ultimately, our work has demonstrated the potential for learning-based approaches to address critical challenges in decentralized multiagent trajectory planning, and we believe that these approaches will play an essential role in enabling the deployment of multiagent systems in real-world applications.

### v4

@@ Abstract
In decentralized multiagent trajectory planners, agents need to communicate and exchange their positions to generate collision-free trajectories. However, due to localization errors/uncertainties, trajectory deconfliction can fail even if trajectories are perfectly shared between agents. To address this issue, we first present PARM and PARM*, perception-aware, decentralized, asynchronous multiagent trajectory planners that enable a team of agents to navigate uncertain environments while deconflicting trajectories and avoiding obstacles using perception information. PARM* differs from PARM as it is less conservative, using more computation to find closer-to-optimal solutions. While these methods achieve state-of-the-art performance, they suffer from high computational costs as they need to solve large optimization problems onboard, making it difficult for agents to replan at high rates. To overcome this challenge, we present our second key contribution, PRIMER, a learning-based planner trained with imitation learning (IL) using PARM* as the expert demonstrator. PRIMER leverages the low computational requirements at deployment of neural networks and achieves a computation speed up to $5614$ times faster than optimization-based approaches.
Supplementary Material
@@ I Introduction
In recent years, multiagent UAV trajectory planning has been extensively studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. In real-world deployments of multiagent trajectory planning methods, it is crucial to deal with challenges such as (1) detecting and avoiding collisions with unknown obstacles, (2) handling localization errors/uncertainties, (3) achieving scalability to a large number of agents, and (4) enabling fast and efficient computation for onboard replanning and quick adaptation to dynamic environments. However, finding effective solutions to these challenges remains an open question.
@@
TABLE I: State-of-the-art UAV Trajectory Planners
Method Multiagent Perception-aware Learning-based
EGO-Swarm [7]
DMPC [23]
MADER [4]
decMPC [24]
Micro Swarm [18] Yes No No
RMADER [5, 6]
MRNAV [16]
DREAM [19]
HDSM [17]
Raptor [25]
Time-opt [26]
PANTHER [27] No Yes No
PA-RHP [28]
PA-RHP [29]
Deep-PANTHER [21]
CGD [30] No Yes Yes
LPA [31]
PUMA [20] No
PARM (proposed) Yes Yes No
PRIMER (proposed) Yes
@@
Many optimization-based approaches [4, 5, 13, 7, 20] have been proposed for multiagent trajectory generation. However, these methods often require significant computational resources, making them challenging to deploy in dynamic environments where fast, on-the-fly replanning is essential. In particular, when optimizing perception-aware trajectories while considering multiagent interactions, computation times increase significantly, as demonstrated in Section III and [20], which hinders real-time deployment. To address this issue, researchers have explored imitation learning (IL)-based approaches [21, 32, 33], which offer the advantage of faster replanning while still producing near-optimal trajectory generation.
@@
    PARM/PARM* – decentralized, asynchronous, perception-aware, multiagent trajectory planner.
@@ II Perception-aware Multiagent Trajectory Generation
This section outlines our perception-aware multiagent planning approach, encompassing both optimization-based and IL-based methods. First, we introduce PARM and PARM*, which are optimization-based perception-aware multiagent trajectory planners. These planners utilize the Robust MADER trajectory deconfliction framework [6] to ensure safe and collision-free trajectories. By considering both position and yaw, PARM and PARM* are capable of tracking multiple obstacles while generating trajectories. Next, we present PRIMER, an IL-based planner that is trained using PARM* as an expert. PRIMER offers fast computation by leveraging the benefits of IL. By incorporating the expertise of PARM*, PRIMER is able to generate effective trajectories while significantly reducing computational requirements. As Table II shows PRIMER is the first learning-based perception-aware multiagent trajectory planner designed to track multiple obstacles while accounting for both position and yaw.
@@ II-A PARM / PARM* —Optimization-based Decentralized, Asynchronous Perception-Aware Multiagent Planning
Our previous work, PANTHER [27], proposed a perception-aware trajectory planner for a single agent in dynamic environments, generating trajectories to avoid obstacles while keeping them in the sensor FOV. In [21], PANTHER* improved the original PANTHER with less conservatism by including separating planes and trajectory total time as optimization variables, but both were limited to tracking and avoiding only one obstacle at a time. To address this limitation, we modify the optimization problem solved by PANTHER to enable the tracking and avoidance of multiple obstacles, resulting in PARM—a decentralized, asynchronous, perception-aware multi-agent trajectory planning system. PARM incorporates this modified optimization approach into the RMADER [6] deconfliction framework, which solves a feasibility linear problem between its own committed trajectories and the shared trajectories from other agents (see Fig. 1 and [4, 6] for details). To facilitate the tracking of multiple obstacles and agents, we modify the FOV term as described in Section IV of [27] as follows.
@@
Additionally, following the method outlined in [21], which minimizes trajectory time in the optimization formulation and consequently reduces trajectory costs, we introduce PARM*, a method that generates closer-to-optimal trajectories. However, this more complex formulation increases computation time, resulting in slower replanning rates.
@@
Note that coupled trajectory generation has the advantage of enabling the optimizer to simultaneously consider both position and yaw trajectories, taking into account the impact of each on the other. On the other hand, the decoupled approach involves optimizing either the position or yaw trajectory first, and then optimizing the other trajectory based on the pre-determined trajectory. While this approach may reduce computation and complexity, it could also result in sub-optimal position and yaw trajectories overall, as the two trajectories are not jointly optimized. Therefore we designed PARM* in a way that they optimize both position and yaw in a coupled manner (See Table II).
@@ II-B PRIMER —Imitation Learning-based PARM*
Similar to PARM/PARM*, PRIMER leverages Robust MADER’s trajectory deconfliction approach, ensuring that all trajectories generated by PRIMER undergo deconfliction checks to guarantee collision-free trajectory generation between agents. Unlike traditional optimization-based methods, which ensure collision-free trajectories by incorporating trajectories from other agents into their constraints, NN-based planners do not inherently provide such guarantees. However, the overall architecture ensures that the final committed trajectories at the conclusion of the replanning steps are collision-free.
@@ III Simulation Results
[…]
@@ III-B Multiagent and multi-obstacle benchmarking
Table IV compares the average performance of PARM, PARM*, and PRIMER with (1) one agent with two obstacles, and (2) three agents with two obstacles. The metrics used to evaluate the performance are as follows:
@@
    Computation time: the time to replan at each step.
@@
    Translational dynamic constraints violation rate: the violation rate of the max velocity, acceleration, and jerk.
@@
    Yaw dynamic constraints violation rate: the violation rate of the maxi yaw rate.
@@
Fig. 5 illustrates (a) Computation Time, (b) Travel Time, and (c-d) Trajectory Smoothness. While the 6-trajectory approach finds smoother trajectories, it requires much longer computation time, resulting in increased travel time. The trajectories generated by PRIMER exhibit smoothness comparable to those of PARM*, while PRIMER have significantly faster computation times compared to PARM and PARM*.
@@ IV Conclusions
To overcome this challenge, we presented PRIMER, a learning-based planner that was trained with imitation learning (IL) using PARM* as the expert demonstrator. PRIMER is a computationally-efficient deep neural network and achieved a computation speedup of up to $5614$ times faster than optimization-based approaches, while maintaining high performance. This speedup enables scalability to a large number of agents, making PRIMER a promising approach for large-scale swarm coordination.
Moving forward, our future work will focus on larger-scale simulations and hardware flight experiments to demonstrate the scalability and performance of PRIMER in complex environments with many agents and obstacles.